Real money balances are defined as M/P. This name should make sense as it is analogous to the real
wage (W/P), real rental price of capital (R/P), and real GDP(PY/P). If the GDP growth rate of a country is
8 percent and inflation is 4 percent, and the velocity of money remains constant, what is the change in
real money balances?
Answer #1

MV=PY (from quantity theory of money)

where, M=money supply, V=velocity of money, P=price level, Y=GDP

Now, Percentage change in M+Percentage change in V=Percentage change in P+Percentage change in Y

or, Percentage change in M = Percentage change in P+Percentage change in Y - Percentage change in V

or, Percentage change in M = 4+8-0 = 12%

Then, percentage change in real money balances(M/P) = Percentage change in M/Percentage change in P = 12%/4% = 3%